in triangle abc, the measure of angle b is three times that of angle a. the measure of angle c is 20 degrees more than that of angle a. find angles a,b,c.
Answer by algebrapro18(249)

Well lets start off by stating that the sum of the angles in a triangle need to add up to 180 degrees. Lets let the measure of angle a = x. Then the measure of angle b = 3x and the measure of angle c = 20+x. Now we just add these and set the sum equal to 180 and solve for x.

x+3x+20+x = 180 combining like terms
5x+20 = 180 subtracting 20 from both sides
5x = 160 dividing both sides by 5
x = 32

So we now know that the measure of angle a is 32 degrees. That means that the measure of angle b is 32*3 = 96 degrees and that the measure of angle c is 20+32 = 52 degrees.